OpenStudy (luigi0210):

Integrate:

OpenStudy (luigi0210):

\[\LARGE \int_{1}^{9}~\frac{3x-2}{\sqrt{x}}~dx\]

OpenStudy (inkyvoyd):

divide into two fractions, and after simplifcation you should get 3sqrt(x)-2(x)^(-1/2)

OpenStudy (inkyvoyd):

then, use the power rule for integrals and evaluate the integral.

OpenStudy (luigi0210):

\[\LARGE \int_{1}^{9}~\frac{3x}{\sqrt{x}}-\int_{1}^{9}\frac{2}{\sqrt{x}}\]

OpenStudy (inkyvoyd):

yes, with the dx's and all. just use laws of exponents (algebra) to help reduce the integrals.

OpenStudy (luigi0210):

Alrighty, thanks inky!